# Schedule On the Fly: Diffusion Time Prediction for Faster and Better Image Generation

### v2

@@ Abstract
Diffusion and flow models have achieved remarkable successes in various applications such as text-to-image generation. However, these models typically rely on the same predetermined denoising schedules during inference for each prompt, which potentially limits the inference efficiency as well as the flexibility when handling different prompts. In this paper, we argue that the optimal noise schedule should adapt to each inference instance, and introduce the Time Prediction Diffusion Model (TPDM) to accomplish this. TPDM employs a plug-and-play Time Prediction Module (TPM) that predicts the next noise level based on current latent features at each denoising step. We train the TPM using reinforcement learning, aiming to maximize a reward that discounts the final image quality by the number of denoising steps. With such an adaptive scheduler, TPDM not only generates high-quality images that are aligned closely with human preferences but also adjusts the number of denoising steps and time on the fly, enhancing both performance and efficiency. We train TPDMs on multiple diffusion model benchmarks. With Stable Diffusion 3 Medium architecture, TPDM achieves an aesthetic score of 5.44 and a human preference score (HPS) of 29.59, while using around 50% fewer denoising steps to achieve better performance. We will release our best model alongside this paper.
footnotetext: The project was initiated and supported by MAPLE Lab at Westlake University.footnotetext: $\dagger$ The work was done during an internship at MAPLE lab.footnotetext: Corresponding author.
@@ 1 Introduction
However, inference with a diffusion model always involves manually selecting a noise scheduler, e.g. how the noise level changes step by step when denoising a clean image from Gaussian noise. This requires the user to delicately adjust parameters like the number of steps, resulting in a high threshold for customers to use. Numerous works attempt to find the optimal scheduler to liberate the users from these tedious tasks.
@@
The leading flow-based models, Stable Diffusion 3 and Flux, provide recommended schedulers that adjust noise levels only based on the targets’ resolution. Sabour [46] and Xia [60] explore ways to fine-tune the schedule for a model, improving either its efficiency or overall performance. In addition, some new one-step generators [16, 33, 34, 35] also achieve impressive performance. Despite the excellent performance they achieve, most of these works hold the assumption that there exists a universally applicable schedule that is optimal for all prompts and images, which is doubtful.
Take several images in Fig. 4 as an example. Images on the right are rich in content, requiring more denoising steps to capture finer details. In contrast, the images on the left are relatively simple and can be generated with fewer steps without compromising quality. Moreover, Karras [20] also proves that different noise schedules also greatly affect the generation quality. Thus, we may ask: Is it possible to adaptively adjust both the number of denoising steps and the noise level at each step during inference, without any manual intervention from the user?
In this paper, we propose Time Prediction Diffusion Models(TPDMs) that can adaptively adjust both the number of steps and the denoising strength during inference. This is achieved by implementing a plug-and-play Time-Prediction Module (TPM) that can predict the next diffusion time conditioned on the latent features at the current step, so that the noise schedule can be adjusted on the fly.
TPM is trained with reinforcement learning. We regard the multi-step denoising process as an entire trajectory, and the quality of the generated image discounted by the number of steps as the reward. The image quality is measured by a reward model aligned with human preference [62].
TPM can be easily integrated into any diffusion models with marginal additional computation, and allow them to automatically adjust hyper-parameters like sample steps and the noise level at each step, reaching a balance between image quality and efficiency without human interventions. Moreover, during training, the model rolls out the diffusion process the same as in inference, directly optimizing the inference performance and reducing denoising steps.
We implement TPDM on several start-of-the-art models, including Stable Diffusion and Flux. With an adaptive noise schedule, our model can also generate $50\%$ fewer steps on average with image quality on par or slightly better (0.322 CLIP-T, 5.445 Aesthetic Score, 22.33 Pick Score, 29.59 HPSv2.1) than Stable Diffusion 3. These results demonstrate that TPDM has the potential to either pursue high-quality image generation or improve model efficiency.
@@
  • We introduce the Time Prediction Diffusion Model (TPDM) that can adaptively adjust the noise schedule during inference, achieving a balance between image quality and model efficiency.
  • To train TPDM, we maximize the generated image quality discounted by the number of steps with reinforcement learning, directly optimizing final performance and efficiency.
  • Our model demonstrates improved performance across multiple evaluation benchmarks, achieving better results with fewer inference steps.
@@ 2.1 Diffusion Models
Diffusion Probability Models (DPMs) [13, 50] recover the original data from pure Gaussian noise by learning the data distribution across varying noise levels. With their strong adaptability to complex data distributions, diffusion models have achieved remarkable performance in diverse fields such as images [39, 40, 42, 15, 21], videos [14, 4, 63], and others [65, 57, 9, 64], significantly advancing the capabilities of Artificial Intelligence Generated Content.
@@ 2.2 Noise Schedule
In order to generate an image, the model must determine the diffusion time for each step. This can be achieved using either discrete-time schedulers [13, 51] or continuous-time schedulers [29, 26] depending on the model. Typically, the diffusion time reflects the noise strength at each step, and most existing approaches rely on pre-determined schedules. Currently, the leading flow-based models, Stable Diffusion 3[10] and FLUX[23], provide recommended schedulers that adjust noise levels only based on the target resolution.
There exist some methods fine-tuning the scheduler to speed up sampling or improve image quality. Xia et al. [60] predict a new denoising time for each step to find a more accurate integration direction, Sabour et al. [46] uses stochastic calculus to find the optimal sampling plan for different schedulers and different models. Wang et al. [56] leverages reinforcement learning to automatically search for an optimal sampling scheduler. Also, adjust the schedule is also effective is also effective in other areas such as molecular generation [17]. Some one-step generators[34, 35, 16] with diffusion distillation [31] also achieve impressive performance.
All of the above noise schedulers not only have high thresholds that require users to adjust many parameters, but also use the same denoising schedule for all prompts and images. On the contrary, TPDM can adaptively adjust the noise schedule during inference, and dynamically select the optimal sampling plan with decreased number of sampling steps for all prompts, achieving a balance between image quality and model efficiency. We will introduce the detailed definitions and practical algorithms of TPDM in Sec. 3.
@@ 2.3 Reinforcement Learning and Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) has recently gained significant attention in the field of large language models (LLMs) [48, 1, 53] and is gradually expanding into other domains. Advances in diffusion models have increasingly incorporated reward models to enhance alignment with human preferences [25, 62, 59]. By treating latents as actions, diffusion models can be optimized with policy gradient (DDPO [3], DPOK[12]), actor-critic framework (DACER [58]), direct preference optimization (Diffusion-DPO [55]) and other reinforcement algorithms. Some recent works have also studied the RLHF for few-step generative models [32, 37]. Since these approaches rely on the Gaussian reverse processes of the diffusion SDE sampler, however, they are challenging to apply in flow models. On the contrary, in this paper we treat the diffusion time as the action instead of latents. Our goal is to adjust the schedule to achieve better quality with fewer steps, offering a more general and flexible solution.
@@ 3 The Proposed Approach
In this section, we first provide a brief review of the fundamental principles of diffusion models, followed by an introduction to the Time Prediction Module (TPM). Finally, we detail the training algorithm for the TPM.
@@ 3.1 Preliminary
Diffusion Models learn to generate images with a reverse process that gradually removes noise from a sample. The leading paradigm for implementing this reverse process is flow matching [27, 10, 11]. Thus we introduce how flow-matching models work, and the detailed structure inside current start-of-the-art models here.
@@
During inference, suppose we generate an image with $N$ steps, each step has a time $t_{n}$ corresponding to its noise strength. Then, the $n$-th generation step can be formulated as:
$x_{t_{n+1}}=x_{t_{n}}+(t_{n+1}-t_{n})\cdot v_{\theta}(x_{t_{n}},t_{n})$ (3)
Usually, in a typical flow matching algorithm, there is a pre-determined schedule that arranges $t_{n}$ every step.
Currently, many state-of-the-art diffusion models are built upon the DiT architecture [24, 8, 10], which only focus on conditional image generation and employ a modulation mechanism to condition the network on both the diffusion time during the diffusion denoising step and the text prompts.
With such a model, we can perform a single denoising step given the latent feature and current noise level.
@@ 3.2 Time Prediction Diffusion Model (TPDM)
As aforementioned, we need a series of denoising steps to generate an image with the trained diffusion model. Usually, a fixed noise schedule is applied in this process for all prompts, assigning a pre-determined noise level for each step.
On the contrary, to enable the model to adjust the noise schedule on the fly, the TPDM predicts how noise strength $t$ decays between adjacent steps. This ensures the noise strength is monotonously decreasing to reflect the denoising progress, avoiding backward progression. To be compatible with training in Sec. 3.3, TPM estimates the distribution over the denoising schedule, rather than predicting an exact value. Suppose we are performing the $n$-th denoising step. Besides the denoising outputs at the current step, TPDM predicts the distribution of the decay rate $r_{n}$ as well. This distribution can be parameterized as a Beta Distribution over $(0,1)$, where the TPDM needs to estimate two parameters $\alpha$ and $\beta$. We notice that when $\alpha>1$ and $\beta>1$ it can ensure a unimodal distribution. Therefore, we reparameterize that TPDM predicts two real numbers $a$ and $b$, and determines the distributions with Eq. 4. Thus the decay rate $r_{n}$ and the next noise level $t_{n}$ can be sampled as in Eq. 5-Eq. 6.
$\alpha=1+e^{a},\qquad\beta=1+e^{b},$ (4)
$r\sim\text{Beta}(\alpha,\beta)$ (5)
@@
TPDM requires only a minimal modification to the original diffusion model: adding a lightweight Time Prediction Module (TPM) as shown in Fig. 3. This module takes the latents before and after the transformer blocks as inputs, so that both the original noisy inputs and the predicted results at this step are taken into consideration. After several convolution layers, TPM pools the latents into a single feature vector and predicts $a$ and $b$ with two linear layers. We also use an adaptive normalization layer [41] in TPM so that the model is aware of the current time embedding.
@@ 3.3 Training Algorithms
To train TPM, we need to roll out at least two denoising steps: predicting the next diffusion time with the first step, and denoising with this time at the next step. A naive method provides a noised image as input to the first step and trains the model with the reconstruction loss calculated at the second step. The gradients would backpropagate through the predicted $t_{n}$ to update TPM in the first step.
However, we found that the trained model tends to complete the denoising process in very few steps during inference, leading to poor image quality. We hypothesize that by supervising the loss calculated after two steps, the model learns to generate a fully denoised image after two steps, and stop intermediately to minimize the loss function. However, what matters is the final image generated after the whole diffusion reverse process, which is ignored in this method.
Therefore, we optimize TPM to maximize the image quality generated after the whole denoising process to achieve precise time prediction. The quality is measured by an image reward model. Considering the computation graph of whole inference is too deep for gradients to backprop, we train the model with Proximal Policy Optimization (PPO) [48], whose loss function is formulated as
@@
Under our problem definition, $s=(c,\epsilon)$ denotes initial states including the input prompt $c$ and gaussian noise $\epsilon\sim\mathcal{N}(0,1)$; $\pi_{\theta}$ denotes the policy network, e.g. our TPDM model. $\pi_{\text{old}}$ denotes the old policy for sampling trajectories; $\pi_{\text{ref}}$ denotes a reference policy for regularization; $y$ denotes the action our policy takes, i.e., the scheduled time; and $\hat{A}(s,y)$ denotes the advantage the action $y$ has given the state $s$.
We will specify the above elements in the PPO including the action and advantage we use in the following.
@@
Usually, when the model makes a sequence of predictions, PPO regards every single prediction as an action and optimizes with them as a batch. Recently, RLOO [1] claims that when the reward signal only appears at the end of the sequence and the environment dynamics is fully deterministic, we can regard the whole sequence as an action with no harm to the performance. Thus for simplicity, the entire generation, including all predicted time in the schedule, is considered as a single action to optimize.
Thus, when computing expectation in Eq. 7, we consider the whole trajectory as a training sample in optimization. However, TPDM only outputs the distribution of each single time prediction, which we denote by $\pi^{(i)}_{\theta}$. By the chain rule, the probability of the entire generation can be calculated as the product of each prediction as in Eq. 8,
$\pi_{\theta}(t_{1},...,t_{N}|s)=\prod_{i=1}^{N}\pi^{(i)}_{\theta}(t_{i}|s,t_{1},...,t_{i-1}),$ (8)
where $N$ denotes the total number of the generation steps. In our model, each factor of the probabilistic policy network $\pi^{(i)}_{\theta}$ is computed through the TPM that outputs a distribution over the possible diffusion time $t_{i}$ for the next step.
@@
When generating samples for policy gradient training, we obtain trajectories by generating images from Gaussian noise with the current TPDM policy. Since there exists no ground truth for these generated images, we choose ImageReward [62] to assign a reward score based solely on the final image in the last time. Considering we do not want a trivially large number of too many steps during generation, we apply a discount factor $\gamma<1$ to discount reward to intermediate diffusion steps and calculate the average over the total number of denoising steps $N$. The final reward for this trajectory is shown in Eq. 9 below,
$R(s,y)=\frac{1}{N}\sum_{k=1}^{N}\gamma^{k-1}\text{IR}(y,s),$ (9)
where IR denotes the image reward model.
We refer readers to RLOO [1] on how to compute the advantage $\hat{A}(s,y)$ directly from a batch $R(s,y)$ without a value model for optimization. This reward function would encourage TPDM to generate images with high quality, and reduce the number of diffusion steps for model efficiency as well. By adjusting $\gamma$, one can control how many diffusion steps tend to be applied to generate high-quality images. The smaller the value of $\gamma$, the fewer steps would tend to be allowed. We will elaborate more in Sec. 4.2.

### v3

@@ Abstract
Diffusion and flow matching models have achieved remarkable success in text-to-image generation. However, these models typically rely on the predetermined denoising schedules for all prompts. The multi-step reverse diffusion process can be regarded as a kind of chain-of-thought for generating high-quality images step by step. Therefore, diffusion models should reason for each instance to adaptively determine the optimal noise schedule, achieving high generation quality with sampling efficiency. In this paper, we introduce the Time Prediction Diffusion Model (TPDM) for this. TPDM employs a plug-and-play Time Prediction Module (TPM) that predicts the next noise level based on current latent features at each denoising step. We train the TPM using reinforcement learning to maximize a reward that encourages high final image quality while penalizing excessive denoising steps. With such an adaptive scheduler, TPDM not only generates high-quality images that are aligned closely with human preferences but also adjusts diffusion time and the number of denoising steps on the fly, enhancing both performance and efficiency. With Stable Diffusion 3 Medium architecture, TPDM achieves an aesthetic score of 5.44 and a human preference score (HPS) of 29.59, while using around 50% fewer denoising steps to achieve better performance.
footnotemark: footnotetext: The project was supported by MAPLE Lab at Westlake University.footnotetext: $\dagger$ The work was done during an internship at MAPLE lab.footnotetext: Corresponding author. G.-J. Qi conceived and formulated the idea.
@@ 1 Introduction
Inference with a diffusion model involves selecting a noise scheduler, e.g. how the noise level changes step by step when denoising from Gaussian noise. This naturally forms a kind of chain-of-thought [59], as it determines how the model gradually generates a real image. Many works reveal that different noise schedules could greatly affect the model performance [19]. The leading flow-matching models, Stable Diffusion 3 [10] and FLUX [22], provide recommended noise schedulers according to the targets’ resolution. Sabour [46] and Xia [61] explore ways to fine-tune the schedule for a model, improving either its efficiency or overall performance. In addition, some one-step generators [16, 33, 34, 35] also achieve impressive performance. Despite the excellent performance they achieve, most of these works hold the assumption that there exists a universally applicable schedule that is optimal for all prompts and images, which is doubtful.
@@
Let us see some examples in Fig. 4. The rightmost images contain more complex visual structures generated from longer prompts. These examples make it necessary for the diffusion model to engage a longer chain of sampling steps to generate richer contents. In contrast, the images on the left have simpler structures (e.g., a single object and uniform background) that can be generated with a shorter chain of fewer steps. Moreover, as discussed later, the noise level set by diffusion time $t$ for each step matters a lot in properly determining how a sample should be denoised at various steps to eventually generate a high quality image. Thus, we ask: Is it possible to adaptively determine both the number of denoising steps and the noise level of each step for each chain of reverse diffusion steps leading to its generated image?
In this paper, we propose Time Prediction Diffusion Models (TPDMs) that can adaptively adjust the noise level at each step as well as the total number of steps during inference. This is achieved by a plug-and-play Time Prediction Module (TPM) that can predict the next diffusion time conditioned on the latent features at the current step. By regarding the reverse diffusion process as a chain of multiple decisions on the diffusion time at each step, we implement reinforcement learning to maximize the reward computed by a reward model aligned with human preference [63]. The reward does not only reflect the quality of the generated image, but is also discounted by the number of steps used in the generation process. Thus, TPDM is directly optimized to generate high-quality images with reduced denoising steps. Moreover, TPM can be easily integrated into any diffusion models with marginal additional computation, allowing them to dynamically adjust the noise schedule for better performance and efficiency.
We implement TPDM on several state-of-the-art models, including Stable Diffusion 3 and Flux. With adaptive noise schedulers, TPDM can generate images with $50\%$ fewer steps on average, while keeping the quality on par or slightly better (0.322 CLIP-T, 5.445 Aesthetic Score, 22.33 Pick Score, 29.59 HPS v2.1) than Stable Diffusion 3. We visualize some generated images in Fig. 1 as well as the number of steps used. These results demonstrate that TPDM has the potential to either pursue high-quality image generation or improve model efficiency.
@@
  • We introduce the Time Prediction Diffusion Model (TPDM). It can predict next diffusion time at each denoising step, determining the optimal noise schedule for each inference instance.
  • We train TPDM with reinforcement learning and maximize a reward with regard to both the quality of the generated image and the number of steps, directly optimizing the final performance and efficiency.
  • TPDM demonstrates improved performance across multiple evaluation benchmarks, achieving better results with fewer inference steps.
@@ 2.1 Diffusion Models
Diffusion Probability Models (DPMs) [13, 50] recover the original data from pure Gaussian noise by learning the data distribution across varying noise levels. With their strong adaptability to complex data distributions, diffusion models have achieved remarkable performance in diverse fields such as images [39, 40, 42, 15, 20], videos [14, 4, 64, 28], and others [66, 57, 9, 65, 27], significantly advancing the capabilities of Artificial Intelligence Generated Content.
@@ 2.2 Noise Schedule
In order to generate an image, the model must determine the diffusion time for each step. This can be achieved using either discrete-time schedulers [13, 51] or continuous-time schedulers [29, 25] depending on the model. Typically, the diffusion time reflects the noise level at each step, and most existing approaches rely on pre-determined schedulers. Currently, the leading flow-matching models, Stable Diffusion 3[10] and FLUX[22], provide recommended schedulers that adjust noise levels only based on the target resolution.
Some methods optimize the scheduler to speed up sampling or improve image quality. Xia et al. [61] predict a new diffusion time for each step to find a more accurate integration direction, Sabour et al. [46] use stochastic calculus to find the optimal sampling plan for different schedulers and different models. Wang et al. [56] leverage reinforcement learning to automatically search for an optimal sampling scheduler. Also, adjusting the scheduler is also effective in other areas such as molecular generation [17]. Some one-step generators[34, 35, 16] with diffusion distillation [31] also achieve impressive performance.
All of the above noise schedulers not only have high thresholds that require users to adjust some hyper-parameters, but also use the same denoising schedule for all prompts and images. On the contrary, TPDM can adaptively adjust the noise schedule during inference and select the optimal sampling plan with suitable sampling steps for each generation, enhancing image quality and model efficiency. We will introduce the detailed definitions and practical algorithms of TPDM in Sec. 3.
@@ 2.3 Reinforcement Learning and Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) has recently gained significant attention in the field of large language models (LLMs) [48, 1, 53] and is gradually expanding into other domains. Advances in diffusion models have increasingly incorporated reward models to enhance alignment with human preferences [23, 63, 60]. By treating latents as actions, diffusion models can be optimized with policy gradient (DDPO [3], DPOK[11]), actor-critic framework (DACER [58]), direct preference optimization (Diffusion-DPO [55]) and other reinforcement algorithms. Some recent works have also studied the RLHF for few-step generative models [32, 37]. Since these approaches rely on the Gaussian reverse processes of the diffusion SDE sampler, however, they are challenging to apply in flow-matching models. On the contrary, in this paper we treat the diffusion time as the action instead of latents. Our goal is to adjust the schedule to achieve better quality with fewer steps, offering a more general and flexible solution.
@@ 3 The Proposed Approach
In this section, we first briefly review the fundamental principles of diffusion models, followed by an introduction to the Time Prediction Diffusion Model (TPDM). Finally, we detail the TPDM’s training algorithm.
@@ 3.1 Preliminary
Diffusion models learn to generate images with a reverse process that gradually removes noise from a sample. The leading paradigm is flow matching [25, 10]. Thus, we introduce how flow-matching models work, and the detailed structure inside current state-of-the-art models here.
@@
During inference, suppose we generate an image with $N$ steps, each step has a time $t_{n}$ corresponding to its noise level. Then, the $n$-th generation step can be formulated as:
$x_{t_{n}}=x_{t_{n-1}}+(t_{n}-t_{n-1})\cdot v_{\theta}(x_{t_{n-1}},t_{n-1})$ (3)
In other words, before the clean image $x_{t_{N}}$ is finally generated, the model forms a chain of $\{x_{t_{n}}\}_{n=1}^{N-1}$ as intermediate results. Usually, in a typical flow matching algorithm, there is a pre-determined schedule for each $t_{n}$, and the proposed TPDM will dynamically decide it for individual samples.
Currently, many state-of-the-art diffusion models are built upon the DiT architecture [22, 8, 10]. They employ multiple layers of transformers to condition the network on both the diffusion time and the text prompts during denoising steps. In this paper, we will build our TPDM on DiT for the sake of a fair comparison.
@@ 3.2 Time Prediction Diffusion Model (TPDM)
To enable the model to adjust the noise schedule on the fly, the TPDM predicts the next diffusion time at each denoising step, as shown in Fig. 2. This can be done by adding a lightweight Time Prediction Module (TPM) to the diffusion model as shown in Fig. 3. This module concatenates the latent features before and after the DiT blocks as inputs, so that both the original noisy inputs and the denoised results at this step are taken into consideration. Then, after several convolution layers, TPM pools the latent features into a single feature vector for prediction. Moreover, we also use an adaptive normalization layer [41] in TPM so that the model is aware of the current diffusion time.
Diffusion time sets the noise level, which should monotonically decrease throughout the denoising process. To prevent backward progression, at each step, TPM predicts decay rate $r$ instead, which quantifies how much diffusion time $t$ decreases between adjacent steps. Since the reinforcement learning-based training algorithm in Sec. 3.3 requires a probability of the predicted time, the final stage of TPM employs two linear layers to process the pooled features and predict the parameters of distribution of $r$ rather than predicting a deterministic value.
Suppose that we are at the $n$-th denoising step. The distribution of the decay rate $r_{n}$ is modeled as a Beta distribution over $(0,1)$ with two parameters $\alpha_{n}$ and $\beta_{n}$ predicted by TPDM. We note that ensuring $\alpha_{n}>1$ and $\beta_{n}>1$ results in a unimodal distribution, which is desired as it avoids TPDM from sampling a potentially vague decay rate from a multi-mode beta distribution. To enforce this constraint, we reparameterize the model such that TPDM predicts two real-valued parameters, $a_{n}$ and $b_{n}$, from which $\alpha_{n}$ and $\beta_{n}$ are determined using Eq. 4. Consequently, the decay rate $r_{n}$ and the next diffusion time $t_{n}$ can be sampled as in Eq. 5 and Eq. 6.
$\alpha_{n}=1+e^{a_{n}},\qquad\beta_{n}=1+e^{b_{n}},$ (4)
$r_{n}\sim\text{Beta}(\alpha_{n},\beta_{n})$ (5)
@@ 3.3 Training Algorithms
To train TPM, we need to roll out at least two denoising steps: predicting the next diffusion time with the first step, and denoising with it at the next step. A naive method provides a noised image as input to the first step and trains the model with the reconstruction loss calculated at the second step. The gradients would backpropagate through the predicted $t_{n}$ to update TPM in the first step. However, we found that the trained model tends to complete the denoising process in very few steps during inference, leading to poor image quality. We hypothesize that by supervising the loss calculated after two steps, the model learns to generate a fully denoised image after two steps, and stop intermediately to minimize the loss function. However, what matters is the final image generated after the whole diffusion reverse process, which is ignored in this method.
After summarizing the above failure, we optimize TPM to maximize the image quality generated after the whole denoising process to achieve precise time prediction. The quality is measured by an image reward model [63]. Considering the computational graph of whole inference is too deep for gradients to backpropagate, so we train the model with Proximal Policy Optimization (PPO) [48], whose loss function is formulated as
@@
Under our task, $s=(c,\epsilon)$ denotes initial states including the input prompt $c$ and gaussian noise $\epsilon\sim\mathcal{N}(0,1)$; $\pi_{\theta}$ denotes the policy network, e.g. our TPDM model. $\pi_{\text{old}}$ denotes the old policy for sampling trajectories; $\pi_{\text{ref}}$ denotes a reference policy for regularization; $y$ denotes the action our policy takes, i.e., the scheduled time; and $\hat{A}(s,y)$ denotes the advantage the action $y$ has given the state $s$. We will specify them in the following.
@@
Usually, when the model makes a sequence of predictions, PPO regards every single prediction as an action and optimizes with them as a batch. Recently, RLOO [1] claims that when the reward only appears at the end of the sequence and the state transition is fully deterministic given a sequence element, we can regard the whole sequence as an action with no harm to the performance. Thus for simplicity, the entire generation, including all predicted time in the schedule, is considered as a single action to optimize.
Thus, when computing expectation in Eq. 7, we consider the whole trajectory as a training sample in optimization. However, TPDM only outputs the distribution of each single time prediction, which we denote by $\pi^{(n)}_{\theta}$ for the $n$-th step. By the chain rule, the probability of the entire generation can be calculated as the product of each prediction:
$\pi_{\theta}(t_{1},...,t_{N}|s)=\prod_{n=1}^{N}\pi^{(n)}_{\theta}(t_{n}|s,t_{1},...,t_{n-1}),$ (8)
where $N$ denotes the total number of the generation steps. In our model, each factor $\pi^{(n)}_{\theta}$ is computed based on the TPM’s output beta distribution at the corresponding step.
@@
When generating samples for policy gradient training, we obtain trajectories by generating images from Gaussian noise with the current TPDM policy. Since there exists no ground truth for these generated images, we choose ImageReward [63] to assign a reward score based solely on the final image in the last denoising step. We aim to enhance final image quality without endlessly increasing denoising steps, so we apply a discount factor $\gamma<1$ (a hyper-parameter) to discount reward to intermediate diffusion steps and calculate the average over the total number of denoising steps $N$. The final reward for this trajectory is shown in Eq. 9 below
$R(s,y)=\frac{1}{N}\sum_{n=1}^{N}\gamma^{N-n}\text{IR}(y,s)$ (9)
where $\text{IR}(y,s)$ denotes the reward of the image generated according to the initial state $s$ and the predicted action $y$. We refer readers to RLOO [1] on how to compute the advantage $\hat{A}(s,y)$ directly from a batch of $R(s,y)$ without a value model for optimization.
This reward function would encourage TPDM to generate images with higher quality and more efficiently. By adjusting $\gamma$, we can control how fast the reward calculated at the final step decays when propagating to previous steps. If a smaller $\gamma$ is set, the reward will decay faster as the number of steps increases. Thus, the model learns to allocate fewer steps in general. We will elaborate more in Sec. 4.2.